OpenStudy (anonymous):

Solve each system by substitution. -4x+3y-2z=7 2x-2y+3z=15 -x+2y-2z=-6 Medal will be given

OpenStudy (anonymous):

Ok -4x+3y-2z=7 3y = 7 + 4x +2z y = 7/3 + 4/3x + 2/3 z (equation 1) substitute y to 2x-2y+3z=15, then 2x-2y+3z=15 2x - 2(7/3 + 4/3x + 2/3 z) + 3z = 15 2x - 14/3 -8/3x - 4/3 z + 3z = 15 2x - 8/3x - 4/3z + 3z = 15 + 14/3 -2/3x + 5/3z = 59/3 both side times 3, then -2x + 5z = 59 2x = 5z - 59 x = 5/2z - 59/2 (equation 2) substitute x to -x+2y-2z=-6, then -x+2y-2z=-6 (5/2z - 59/2) + 2y - 2z = -6 5/2z - 59/2 + 2y - 2z = -6 5/2z + 2y - 2z = -6+ 59/2 5/2z + 2y - 2z = 47/2 5/2z -2z+ 2y = 47/2 z/2 + 2y = 47/2 both side times 2, then z + 4y = 47 z = 47 - 4y (equation 3) substitute this z (equation 3) to x (equation 2), then x = 5/2z - 59/2 x = 5/2 (47 - 4y ) - 59/2 x = 235/2 - 10y - 59/2 x = 176/2 - 10y (equation 4) OF COURSE substitute this x (equation 4) and z (equation 3) to y (equation 1), then y = 7/3 + 4/3x + 2/3 z y = 7/3 + 4/3 (176/2 - 10y) + 2/3 (47 - 4y) y = 7/3 + 352/3 - 40/3y + 94/3 - 8/3y y + 40/3y + 8/3y = 7/3 + 352/3 + 94/3 y (1 + 40/3 + 8/3) = 7/3 + 352/3 + 94/3 both side times 3 y (3 + 40 + 8) = 7 + 352 + 94 y (51) = 453 y = 453 /51

OpenStudy (anonymous):

from equation 4 we get x = 176/2 - 10y x = 176/2 - 10 (453 /51) x = 176/2 - 4530 /51

OpenStudy (anonymous):

from equation 3 we get z = 47 - 4y z = 47 - 4(453 /51) z = 47 - 1812/51

OpenStudy (anonymous):

Thanks helped a lot

OpenStudy (anonymous):

you are welcome :)